OpenStudy (anonymous):

2. Concrete can be purchased by the cubic yard. How much will it cost to pour a slab 11 feet by 11 feet by 3 inches for a patio if the concrete costs $63.00 per cubic yard? * $211.75 * $70.58 * $317.63 * $1,905.75

OpenStudy (anonymous):

So you'll need to find the volume of the slab to be poured using the formula for volume of a rectangle. Then that's your cubic yardage needed, mulitply that by the price for concrete. Where are you having trouble?

OpenStudy (anonymous):

i just dont understand how to find the volume after that ill get it

OpenStudy (anonymous):

Volume formula is Length times Width time Height. But you have to make sure that all three dimensions are in the unit asked for, so in terms of yard. So start by converting all the given dimensions to yard, then find volume using the formula

OpenStudy (anonymous):

so 132*132*3= 52,272 then convert to yards?

OpenStudy (anonymous):

so 1452* 63.00?

OpenStudy (anonymous):

Hang on, I've gotta make sense of it and catch up now, because I would have converted everything to yards from the start, rather than converting down to inches and then back up.

OpenStudy (anonymous):

mk =)

OpenStudy (anonymous):

Yes, you were right. 1452 * 63 should give you the correct answer.

OpenStudy (anonymous):

91,476 isnt a choice?

OpenStudy (anonymous):

Alright, give me a second to do it the way I would have from the start and let me see if I get a different answer

OpenStudy (anonymous):

thanks =)

OpenStudy (anonymous):

Yea. I get one of those choices doing it my way. Do the conversion up to feet first, then find the volume.

OpenStudy (anonymous):

11*11*.25=30.25*63.00=1,905.75!!!!!!!!!!!!!!!!! yayyyyyy thank you so much!

OpenStudy (anonymous):

NO. You only converted to feet. then multiplied by the price per yard. You have to convert all the dimensions up to yards. I'm so sorry I messed that up when I said it.

OpenStudy (anonymous):

I feel pretty bad about that slip @TristenShae , I hope you see this and understand it / correct it before turning that answer in.

OpenStudy (anonymous):

what was the answer???